Dunia Construction Co. (DCC) is considering a new inventory system that will cost
RM750,000. The system is expected to generate positive cash flows over the next
four years in the amounts of RM350,000 in year one, RM325,000 in year two,
RM150,000 in year three, and RM180,000 in year four. DCC's required rate of return
is 8%.
i. What is the net present value of this project?
ii. What is the internal rate of return of this project?
iii. What is the modified internal rate of return of this project?
1
Expert's answer
2015-03-16T10:31:09-0400
Cost RM750,000. RM350,000 in year one, RM325,000 in year two, RM150,000 in year three, and RM180,000 in year four. DCC's required rate of return is 8%.
i. The net present value of this project is:
NPV = -750,000 + 350,000/1.08 + 325,000/1.08^2 + 150,000/1.08^3 + 180,000/1.08^4 = 104,089.4
ii. The internal rate of return of this project is:
IRR = 15.13% - is the rate of return, for which NPV = 0.
iii. The modified internal rate of return of this project is:
MIRR = ((350,000 + 325,000 + 150,000 + 180,000)/750,000)^0.25 - 1 = 7.59%